# AReaL: A Large-Scale Asynchronous Reinforcement Learning System for Language Reasoning

### v4

@@ 3. Answer: [Yes]
  4. Justification: The main scope is in Section˜2 and the main contribution is in Section˜4. The abstract summarizes them.
@@ 13. Answer: [Yes]
  14. Justification: All assumptions have been illustrated in Proposition˜1 and Section˜2
@@ 18. Answer: [Yes]
  19. Justification: We have discussed the model, dataset, and hyper-parameters we use in Appendix˜B and Section˜6.
@@ 38. Answer: [Yes]
  39. Justification: In Appendix˜B and Section˜6.
@@ B.2 Dataset Details
For the math task, we used the open-source data from DeepScaleR [deepscaler2025], For code training, we used the dataset released by DeepCoder [deepcoder2025]. All compared methods use the same dataset.
@@ B.4 Baselines
In our experiments, we use the lastest version (main branch of verl repository, May 7, 2025) of verl [verl-hybridflow] to evaluate the training throughput in Figure˜4 and the training hours in Table˜1. For most of the results, we use SGLang [sglang] v0.4.6 as generation backend and pytorch FSDP [fsdp] as training backend. In a few cases where SGLang raises errors (experiments with 32B models or 64 nodes), we use vLLM [vllm] v0.8.4 as a substitution.

### v5

@@ 3. Answer: [Yes]
  4. Justification: The main scope is in Section˜3 and the main contribution is in Section˜5. The abstract summarizes them.
@@ 13. Answer: [Yes]
  14. Justification: All assumptions have been illustrated in Proposition˜1 and Section˜3
@@ 18. Answer: [Yes]
  19. Justification: We have discussed the model, dataset, and hyper-parameters we use in Appendix˜B and Section˜7.
@@ 38. Answer: [Yes]
  39. Justification: In Appendix˜B and Section˜7.
@@ B.2 Dataset Details
For the math task, we used the open-source data from DeepScaleR [25], For code training, we used the dataset released by DeepCoder [24]. All compared methods use the same dataset.
@@ B.4 Baselines
In our experiments, we use the lastest version (main branch of verl repository, May 7, 2025) of verl [45] to evaluate the training throughput in Figure˜4 and the training hours in Table˜1. For most of the results, we use SGLang [63] v0.4.6 as generation backend and pytorch FSDP [62] as training backend. In a few cases where SGLang raises errors (experiments with 32B models or 64 nodes), we use vLLM [18] v0.8.4 as a substitution.